Previous

Calendar

February

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD27

28

JD28

29

JD29 (-20 C, peak gusts 25 m/s)
Site Visit (JB, SpB): Details , Photos (11)

re-deployed the re-calibrated CSAT3 sonic (sn 0226) at 2.5 m on the mini tower, removed the 2.5 m CSAT sonic (sn 0438) to send back for re-cal

JB and SpB helped Duane load NT's snow blower onto the snow cat

SpB flagged off an area containing subterranean sensors around the mini tower...fortify this with rope/string next visit

30

JD30

31

JD31

1

JD32
Site Visit (JB, SLS, NT): Details

switched the SDM setting on the re-calibrated sonic from 3 to 4; this got it working fine

reinforced the flagging around the mini tower

power cylcled the gmp343 probe and got it back on scale in cockpit

span @ 200 psi; N2 @ 190 psi; retired the N2 (CC198536 @ 190 psi, started N2 CC125555 @ 400 psi -- this was Jia's summer pressure bomb cylinder)

see details for snow depths

SpB Off

2

JD33

3

JD34

4

JD35

5

JD36

6

JD37

7

JD38
Site Visit (JB): Details

N2 @ 200 psi, span @ 100 psi

swapped 4 flasks at west robot

downloaded from Noah's logger

set up one of Noah's snow sensors (next to a tree bole, "beneath-canopy", near Jia's sap flow site ~30 yards SW of the north canopy tower; not sure if it is working properly)

SpB returns

8

JD39

9

JD40

10

JD41

11

JD42

12

JD43

13

JD44 (beautiful, warm day, though gusty at times)
Site Visit (JB): Details

retired N2 cylinder CC125555 (~empty), started N2 cylinder CC193045 (1900 psi)

retired span cylinder CC18996 (~empty), started span cylinder CC199497 (~1775 psi)

swapped pcmcia cards at the mrs logger

downloaded from the soil, tcs, and noah loggers

connected snow sensor #3

hauled a suitcase full of flasks back to campus to mail to Utah

brought the pressure bomb's regulator, mag glass and stopper kit back to campus for Mick

see details for snow depth and leak check info

14

JD45

15

JD46

16

JD47

17

JD48

18

JD49
at 18:40 MST, UCB logger (chan 202) stops output 1-hz data.

19

JD50

20

JD51 (very warm, sunny, calm)
Site Visit (JB, NT): Details

span @ 1750 psi (dropping 25 psi/week, seems good); N2 @ 1700 psi (dropping ~200 psi/week, seems okay)

downloaded from the UC logger

tried to diagnose the problem with channel 202/ucb logger

see details for info on UCB logger

21

JD52 (cold and snowing)
Site Visit (JB, SpB): Details , Photos (9)

Details

swapped 4 flasks

downloaded from noah's logger

swapped second re-calibrated CSAT3 sonic (sn 0438) with the EOL sonic (sn 0536) @ top of tower (SpB brought the EOL sonic back to NCAR)

adjusted span gas flow rate

swapped the broken ucb logger (cr23x micrologger) with one of Jia's cr10x loggers

see details about this logger swap

22

JD53 (warm, sunny, mostly calm)
Site Visit (JB, JH, NT, SLS): Details

tested the cr23x (ucb) in the trailer and confirmed that it is unresponsive

replaced the entire battery/charger system for the ucb logger

got the data from ucb (Jia's cr10x) going back into the data system again (was a matter of sufficient power to the logger)

started running "ucb10x_2.csi" at the ucb logger (this program uses instruction P138 for the cs616 water content reflectometer), although there still seems to be something going on with this sensor; investigate when the cr23x is returned

23

JD54

24

JD55

25

JD56

26

JD57

27

JD58

28

JD59

29

JD60
Leap day

1

JD61